OpenStudy (anonymous):

Find the limit of the function algebraically. limit as x approaches zero of quantity x squared plus two x divided by x to the fourth power.

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\frac{ x ^{2} + 2x }{ x ^{4} }\]

jimthompson5910 (jim_thompson5910):

it might help to look at a table of values. Specifically x values that are very close to 0 things like x = -0.5 x = -0.4 x = -0.3 x = -0.2 x = -0.1 x = 0 x = 0.1 x = 0.2 x = 0.3 x = 0.4 x = 0.5

OpenStudy (fortytherapper):

I'm assuming L'Hopitals wouldn't be valid in this situation?

jimthompson5910 (jim_thompson5910):

You could use L'Hospitals' rule but only on that current form if you were to simplify like so \[\Large \frac{ x ^{2} + 2x }{ x ^{4} }\] \[\Large \frac{ x(x + 2) }{ x ^{4} }\] \[\Large \frac{ x(x + 2) }{ x*x ^{3} }\] \[\Large \frac{ {\color{red}{x}}(x + 2) }{ {\color{red}{x}}*x ^{3} }\] \[\Large \frac{ {\color{red}{\cancel{x}}}(x + 2) }{ {\color{red}{\cancel{x}}}*x ^{3} }\] \[\Large \frac{ x+2 }{ x ^{3} }\] then L'Hospitals' rule wouldn't work

jimthompson5910 (jim_thompson5910):

In my opinion, it's always a good idea to simplify as much as possible before doing the limit

OpenStudy (fortytherapper):

I agree 100%, I just wasn't sure if L'Hopitals was considered algebraically